the shorter diagonal of rhombus is equal to one of its sides.
(a) find each diagonal of the rhombus.

Answer by venugopalramana(3286)

IN A RHOMBUS IF SIDE IS X WE HAVE
(D1/2)^2+(D2/2)^2=X^2...WHERE D1 AND D2 ARE DIAGONALS
IF D1=X THEN
(X/2)^2+(D2/2)^2=X^2...
(D2/2)^2=3X^2/4
D2/2=X*SQRT3/2
D2=X*SQRT3
SO EACH DIAGONAL IS X AND X*SQRT(3)